# Code-Embed: A Family of Open Large Language Models for Code Embedding

### v2

@@ 3.1 General Training Evaluation
Model Apps CosQA Text2SQL CSN CSN-CCR CodeTrans StackOverFlow CodeFeedBack Avg
@@
Table 1: Performance of the Pretrained CodeXEmbed Model Family across various tasks, including Text-to-Code, Code-to-Text, Code-to-Code, and Hybrid Code, along with their average scores. CSN stands for CodeSearchNet.
@@ 3.3 Unified Text and Code Retrieval
Model BM25 gte-large gte-Qwen2 E5-Mistral CodeXEmbed CodeXEmbed CodeXEmbed
@@
In the 400M models, CodeXEmbed achieves a 0.43 performance boost compared to GTE-large (Li et al., 2023b), on which it is trained. This underscores the advantage of our approach, demonstrating the potential to improve text retrieval performance by incorporating code data. There are not many 2B-sized language models available, and we selected Gemma 2B Team et al. (2024) due to its strong performance in code retrieval. For text retrieval, it achieves comparable performance to gte-Qwen2 of similar size.
Model NV-Embed-V2 SFR-v2 CodeXEmbed
@@
We present the performance of leading text retrieval models from the MTEB leaderboard [^4] https://huggingface.co/spaces/mteb/leaderboard on both code and text tasks, in Table 3. Compared to the top-ranked model, NV-Embed-V2 Moreira et al. (2024) [^5] Top-performing model as of 10/15/2024., our model outperforms it by 4.38 points, with an average score across text and code retrieval datasets.
@@ 3.4.1 RepoEval
For code generation, we supply the top-5 retrieved code snippets as input to the GPT-3.5-turbo generation model. As shown in Table 4, all sizes of CodeXEmbed outperform the canonical setup. While some files may not contain direct solutions, as in the canonical documents, they often include valuable function definitions or usage examples that improve code generation outcomes. This suggests that our embeddings effectively capture the repository structure and retrieve contexts that implicitly support problem-solving.
Model Repo-Level
RepoEval SWE-Bench-Lite
None 23.9 0.7
BM25 30.8 1.0
Voyage, code 43.2 0.7
OpenAI, small 48.0 0.3
OpenAI, rerank 49.6 0.0
CodeXEmbed400M 52.5 0.6
CodeXEmbed2B 66.3 1.6
CodeXEmbed7B 57.8 2.6
Gold 39.1 2.7
Table 4: Performance of repository-level code retrieval augmented generation using gpt-3.5-turbo.
@@ 3.4.2 SWE-Bench-Lite
In our experiments, we use SWE-bench-Lite [^6] https://huggingface.co/datasets/princeton-nlp/SWE-bench_Lite, a curated subset of 300 problems from the original SWE-bench benchmark. It focuses on resolving GitHub issues by requiring models to modify multiple files to pass test cases, providing a manageable and reproducible dataset. SWE-bench-Lite also includes a pre-configured Docker container, ensuring consistent evaluation across systems and further standardizing the testing environment. Our results show that incorporating improved retrieval methods can significantly enhance the end-to-end performance of code retrieval-augmented generation, bringing it closer to using gold content and boosting problem-solving efficiency and accuracy.
@@ 3.5 Impact of the Base Models
To understand the base model’s impact, we examine: (1) if training from a text retrieval model outperforms a generation model, and (2) if a code-specific generation model offers more advantages than a general language model.
3.5.1 Embedding Models v.s. LLMs
@@
Model gte-Qwen2 Gemma-v2 SFR-v2 Mistral
@@
3.5.2 Code-Specific LLMs v.s. General LLMs
@@
Model StarCoder-v2 DeepSeek-Coder Gemma-v2 Mistral
@@
3.6 Programming Language Transferability
Model Size Python Java Go PHP Javascript Ruby SQL AVG $\Delta$-All
All 400M 59.70 72.09 79.28 70.96 70.18 73.01 58.87 69.16 -
2B 65.73 79.40 81.01 79.39 76.76 78.42 58.23 74.14 -
7B 67.56 81.36 85.76 81.89 78.37 83.11 65.74 77.68 -
Python 400M 55.43 49.57 49.56 40.57 44.57 44.57 44.21 46.93 $-32.1\%$
2B 60.86 71.46 68.69 57.88 62.95 72.21 56.09 64.31 $-13.3\%$
7B 67.02 78.35 81.56 70.88 75.73 79.19 62.56 73.61 $-5.2\%$
Java 400M 54.79 78.37 72.94 70.85 69.85 68.78 50.31 66.56 $-3.8\%$
2B 63.03 82.17 79.35 80.30 76.00 76.76 58.03 73.66 $-0.6\%$
7B 66.24 84.25 83.72 80.18 79.82 82.82 64.59 77.37 $-0.4\%$
Table 7: Python/Java indicates that only the Python and Java portions were used to train CodeXEmbed, while All indicates that all programming languages were used to train CodeXEmbed. $\Delta$-All represents the difference between the Python/Java AVG score and the All AVG score for the same model size.
4 Related Work
4.1 Retrieval Models for Text and Code
Text retrieval models have significantly advanced by exploiting supervision from natural language inference tasks and labeled query-document pairs, such as the MS-MARCO passage ranking dataset Bajaj et al. (2016), to train effective text embeddings Izacard et al. (2021); Wang et al. (2022); Xiao et al. (2024). Recently, researchers have leveraged large language models (LLMs) Jiang et al. (2023) as the base for training retrieval models, resulting in state-of-the-art performance. Notable examples include E5-Mistral Wang et al. (2023), SFR-Embedding Meng et al. (2024), and NV-Embedding Lee et al. (2024). While numerous models have been developed for text retrieval tasks, few have focused specifically on code retrieval. Among the few are the Voyage code model AI (2024) and OpenAI’s embeddings Neelakantan et al. (2022); however, both are closed-source models, limiting their accessibility and adaptability for the wider research community.
4.2 Code Retrieval Augmented Generation
Neural code generation has been an important task Lu et al. , and increasingly strong code language models have been developed Roziere et al. (2023); Nijkamp et al. (2023); Li et al. (2023a); Guo et al. (2024); Team (2024) to solve various tasks Chen et al. (2021); Lai et al. (2023); Jimenez et al. (2024). However, most LMs generate code solely from natural language problem descriptions and the models’ parametric knowledge, without leveraging external programming resources or using a retrieval-augmented generation approach. While prior work has focused on text-centric tasks with general-domain corpora like Wikipedia Asai et al. (2024), some studies have used retrieved programming context from repositories Ding et al. (2024); Yang et al. (2024) or documentation Zhou et al. (2023). Code retrieval is crucial for enhancing code generation in RAG systems because it allows models to access relevant external code resources, leading to more accurate and context-aware code outputs. Our code retrieval model demonstrates significant improvements in code RAG performance, underscoring the importance of effective code retrieval in code generation tasks.
5 Conclusion
[…]
6 Limitations
While CodeXEmbed achieves state-of-the-art results in code and text retrieval, there are certain limitations to consider. First, the model’s large size, particularly the 7B parameter variant, leads to significant storage and computational costs. With embeddings of 4096 dimensions, the model is more resource-intensive compared to smaller or more efficient representations. Although embedding size reduction techniques, such as Matryoshka representation learning Kusupati et al. (2022), have shown promise in maintaining performance with lower dimensions, further research is needed to apply these approaches effectively to code retrieval without compromising performance. Additionally, while CodeXEmbed generalizes well across different programming languages and text retrieval tasks, it may still struggle with niche programming domains or highly specialized languages that were underrepresented in the training data. Future work could explore adapting the model for these underrepresented languages and improving efficiency without sacrificing accuracy.
Appendix A Appendix
A.1 Dataset Details
[…]
A.2 Implementation Details
We summarize the hyperparameters in Table 8. For the code training data, we prepend a prompt to the query in the format: “Instruct: Given Code or Text, retrieve relevant content. Query:”.

### v3

@@ 3.1 General Training Evaluation
Model Apps CosQA
Text2SQL
CSN
CSN-CCR
CodeTrans StackOverFlow CodeFeedBack Avg
@@
Table 1: Performance of the CodeXEmbed model family across tasks, along with average scores. CSN refers to CodeSearchNet. Baseline numbers are from the CoIR Leaderboard.
@@ 3.3 Unified Text and Code Retrieval
Dataset BM25 gte-large gte-Qwen2 E5-Mistral CodeXEmbed CodeXEmbed CodeXEmbed
@@
In the 400M models, CodeXEmbed achieves a 0.43 performance boost over GTE-large (Li et al., 2023b), the model it is trained on. This highlights the advantage of our approach, showing the potential to improve text retrieval by incorporating code data. Few 2B-sized language models are available; we selected Gemma 2B Team et al. (2024) for its strong code retrieval performance. For text retrieval, it performs comparably to gte-Qwen2 of similar size.
Dataset NV-Embed-V2 SFR-v2 CodeXEmbed
@@
We present the performance of top text retrieval models from the MTEB leaderboard [^7] https://huggingface.co/spaces/mteb/leaderboard on both code and text tasks, in Table 3. Compared to the top-ranked model, NV-Embed-V2 Moreira et al. (2024) [^8] Top-performing open model as of 3/28/2025., our model surpasses it by 4.38 points, with an average score across text and code datasets.
3.4 Impact of Multi-Stage Training
To evaluate the effectiveness of our multi-stage training approach, we conduct an ablation study on four settings:
  • Single-stage (Baseline): Trained in a single phase using a mixture of text and code retrieval data with a single LoRA adapter.
  • Two-stage (Same LoRA): The first stage is trained on text retrieval; the second stage continues with mixed text and code retrieval, reusing the same LoRA adapter across both stages.
  • Two-stage (Different Optimizers): Uses different LoRA adapters for each stage, with separate optimizers for each.
  • Proposed Two-stage (Separate LoRA): Applies different LoRA adapters for each stage, using a continuous optimizer throughout training.
Training Setup BEIR CoIR
Single-stage (Baseline) 65.79 64.56
Two-stage (Same LoRA) 67.54 58.61
Two-stage (Diff. Optimizer) 66.95 66.06
Proposed Two-stage (Separate LoRA) 67.88 67.41
Table 4: Comparison of training methods on BEIR dev and CoIR test sets (NDCG@10).
Table 4 shows our Proposed Two-stage (Separate LoRA) achieves the best performance on BEIR (67.88) and CoIR (67.41), highlighting the importance of modular adaptation. The Single-stage (Baseline) struggles on CoIR, while Two-stage (Same LoRA) improves BEIR but drops on CoIR. Two-stage (Different Optimizers) balances performance but falls short of our approach.
@@ 3.5.1 RepoEval
For code generation, we supply the top-5 retrieved code snippets as input to the GPT-3.5-turbo generation model. All generator parameters and decoding hyperparameters are kept frozen throughout the experiments. As shown in Table 5, all sizes of CodeXEmbed outperform the canonical setup. While some files may not contain direct solutions, as in the canonical documents, they often include valuable function definitions or usage examples that improve code generation outcomes. This suggests that our embeddings effectively capture the repository structure and retrieve contexts that implicitly support problem-solving.
Dataset None BM25 Voyage OpenAI OpenAI CodeXEmbed Gold
rerank 400M 2B 7B
gpt-3.5
RepoEval 23.9 30.8 43.2 48.0 49.6 52.5 66.3 63.8 39.1
SWE-Bench-Lite 0.7 1.0 0.7 0.3 0.0 0.7 2.0 3.0 2.7
gpt-4o
SWE-Bench-Lite 2.3 - - - 21.7 19.7 21.7 25.0 30.7
Table 5: Performance of repository-level code retrieval augmented generation using gpt-3.5-turbo-0125 and gpt-4o-2024-08-06. CodeXEmbed variants denote models of size 400M, 2B, and 7B.
@@ 3.5.2 SWE-Bench-Lite
In our experiments, we use SWE-bench-Lite [^9] https://huggingface.co/datasets/princeton-nlp/SWE-bench_Lite, a curated subset of 300 problems from the original SWE-bench benchmark. It focuses on resolving GitHub issues by requiring models to modify multiple files to pass test cases, offering a manageable and reproducible dataset. SWE-bench-Lite also includes a pre-configured Docker container, ensuring consistent evaluation across systems and further standardizing the testing environment.
We evaluate our retrieval-augmented generation pipeline using GPT-3.5-turbo as the code generator, with all generator parameters and decoding hyperparameters kept frozen. We also conduct parallel experiments with GPT-4o and observe similar trends, confirming that the performance gains from improved retrieval are consistent across model backends.
As shown in Table 5, our results show that incorporating improved retrieval methods significantly enhances end-to-end performance of code retrieval-augmented generation, bringing it closer to using gold content and boosting problem-solving efficiency and accuracy.
@@ 3.6 Impact of the Base Models
To understand the base model’s impact, we examine: (1) if training from a text retrieval model outperforms a generation model, and (2) if a code-specific generation model offers more advantages than a general language model. We present the ablation study comparing an embedding model to a generation LLM in Appendix A.2.1 and the study on code-specific vs. general LLMs in Appendix A.2.2.
3.7 Programming Language Transferability
Model Size Python Java Go PHP Javascript Ruby SQL AVG $\Delta$-All
All 400M 59.70 72.09 79.28 70.96 70.18 73.01 58.87 69.16 -
2B 65.73 79.40 81.01 79.39 76.76 78.42 58.23 74.14 -
7B 67.56 81.36 85.76 81.89 78.37 83.11 65.74 77.68 -
Python 400M 55.43 49.57 49.56 40.57 44.57 44.57 44.21 46.93 $32.1\%$
2B 60.86 71.46 68.69 57.88 62.95 72.21 56.09 64.31 $13.3\%$
7B 67.02 78.35 81.56 70.88 75.73 79.19 62.56 73.61 $5.2\%$
Java 400M 54.79 78.37 72.94 70.85 69.85 68.78 50.31 66.56 $3.8\%$
2B 63.03 82.17 79.35 80.30 76.00 76.76 58.03 73.66 $0.6\%$
7B 66.24 84.25 83.72 80.18 79.82 82.82 64.59 77.37 $0.4\%$
Table 6: Python/Java indicates that only the Python and Java portions were used to train CodeXEmbed, while All indicates that all programming languages were used. $\Delta$-All represents the difference between the Python/Java AVG score and the All AVG score for the same model size.
To better evaluate retrieval performance across languages, we organize the CoIR test set into language-specific experiments by grouping all subsets associated with the same programming language. No additional resplitting is applied. For each language, we include all CoIR test subsets labeled with that language. The details are shown in Table 7.
Language CoIR Test Subsets Used
Python APPS · CoSQA · CodeSearchNet-Python · CodeSearchNet-CCR-Python · CodeTrans-DL
Java CodeSearchNet-Java · CodeSearchNet-CCR-Java
Go CodeSearchNet-Go · CodeSearchNet-CCR-Go
PHP CodeSearchNet-PHP · CodeSearchNet-CCR-PHP
JavaScript CodeSearchNet-JavaScript · CodeSearchNet-CCR-JavaScript
Ruby CodeSearchNet-Ruby · CodeSearchNet-CCR-Ruby
SQL Synthetic-Text2SQL
Table 7: CoIR test subsets used for each programming language.
4 Related Work
4.1 Retrieval Models for Text and Code
Text retrieval models have significantly advanced by exploiting supervision from natural language inference tasks and labeled query-document pairs, such as the MS-MARCO passage ranking dataset Bajaj et al. (2016), to train effective text embeddings Izacard et al. (2021); Wang et al. (2022); Xiao et al. (2024). Recently, researchers have leveraged large language models (LLMs) Jiang et al. (2023) as the base for training retrieval models, resulting in state-of-the-art performance. Notable examples include E5-Mistral Wang et al. (2023), SFR-Embedding Meng et al. (2024); Meng* et al. (2024), and NV-Embedding Lee et al. (2024). While numerous models have been developed for text retrieval tasks, few have focused specifically on code retrieval Zhang et al. (2024); Suresh et al. (2024). Among the few are the Voyage code model AI (2024), OpenAI’s embeddings Neelakantan et al. (2022) and embeddings for code issue localization Reddy et al. (2025). however, both are closed-source models, limiting their accessibility and adaptability for the wider research community.
4.2 Code Retrieval Augmented Generation
Neural code generation has been an important task Lu et al. (2021), and increasingly strong code language models have been developed Roziere et al. (2023); Nijkamp et al. (2023); Li et al. (2023a); Guo et al. (2024); Team (2024) to solve various tasks Chen et al. (2021); Lai et al. (2023); Jimenez et al. (2024). However, most LMs generate code solely from natural language problem descriptions and the models’ parametric knowledge, without leveraging external programming resources or using a retrieval-augmented generation approach. While prior work has focused on text-centric tasks with general-domain corpora like Wikipedia Asai et al. (2024), some studies have used retrieved programming context from repositories Ding et al. (2024); Yang et al. (2024) or documentation Zhou et al. (2023). Code retrieval is crucial for enhancing code generation in RAG systems because it allows models to access relevant external code resources, leading to more accurate and context-aware code outputs. Our code retrieval model demonstrates significant improvements in code RAG performance, underscoring the importance of effective code retrieval in code generation tasks.
5 Conclusion
In the underexplored field of code retrieval, we present CodeXEmbed, a family of code embedding models ranging from 400M to 7B parameters. Our unified training pipeline integrates multiple programming languages and reformulates diverse code-related tasks within a common retrieval framework. CodeXEmbed ranks as the top model on the CoIR benchmark and achieves performance comparable to SOTA text retrieval models on BEIR. By enhancing retrieval capabilities, we significantly improve end-to-end retrieval-augmented generation for code-related tasks. Bridging the gap between text and code retrieval, we release our models to foster research and innovation in developer tools and programming language understanding.
Appendix A Appendix
A.1 Dataset Details
[…]
Text-to-Code Retrieval: Text2SQL: We follow Text2SQL works in Yu et al. (2018) and Zhong et al. (2017), converting the text part into the queries to retrieval SQL as the documents, derived from Wikipedia data. In contrast, the CoIR benchmark uses synthetic-text2sql, generated using weak and strong LLMs. Code Contest: We extract the Code Contest tasks in the Khan et al. (2024), which has 11 programming language like java, c, c# et al. While CoIR benchmark use APPS, which is the interview question only in python programming language.
Code-to-Text Retrieval: Code Summary: We follow Hu et al. (2018) and Liu et al. (2020), which summarize code based on associated documentation, while CoIR utilizes a CodeSearchNet variant with code-comment pairs for the summary task. Code Clone Detection: We follow Svajlenko & Roy (2015) and Lu et al. (2021) to convert the code as query and its clone code as the docs ; CoIR has no dataset for this task.
Code-to-Code Retrieval: Code Translation: We follows Khan et al. (2024) and Lu et al. (2021), converting code in one language into queries and corresponding code in another language into documents. CoIR has no dataset for this task. Code Completion: We use the code completion tasks from Khan et al. (2024) and Lu et al. (2021), while CoIR uses a modified CodeSearchNet, splitting code snippets into query-document pairs.
Hybrid Code Retrieval: Code Issue Fix: We follow Khan et al. (2024) and Lu et al. (2021) to involve the code issue fix task; CoIR has no dataset for this task. Code Agent Conversation: We process the conversation data from Team et al. , filtering answers based on favorite scores and excluding data used in CoIR.
A.2 Impact of the Base Models
We analyze the base model’s impact by comparing (1) text retrieval vs. generation models for training and (2) a code-specific generation model offers advantages over a general language model.
A.2.1 Embedding Models v.s. LLMs
@@ A.2 Impact of the Base Models
Dataset gte-Qwen2 Gemma-v2 SFR-v2 Mistral
@@
A.2.2 Code-Specific LLMs v.s. General LLMs
@@
Dataset StarCoder-v2 DeepSeek-Coder Gemma-v2 Mistral
@@
A.3 Implementation Details
We summarize the base model detail in Table 10 and hyperparameters in Table 11. For the code training data, we prepend a prompt to the query in the format: “Instruct: Given Code or Text, retrieve relevant content. Query:”. We use
Model Name Model Size Version Number Date of Release
gte-large-en-v1.5 400M v1.5 Approximately November 2023
gemma-2-2b-it 2B v1 Approximately June 2024
Mistral-7B-Instruct-v0.3 7B v0.3 May 2024
Table 10: Model details including name, size, version, and release date.